40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 10 01:09:20 2015, rana, Update, LSC, Moved LSC sensing matrix notch frequencies NewLSCnotches.png
    Reply  Thu Sep 10 04:23:00 2015, ericq, Update, LSC, Moved LSC sensing matrix notch frequencies DRMI_1F.pdfDRMI_3F.pdf
Message ID: 11588     Entry time: Thu Sep 10 01:09:20 2015     Reply to this: 11589
Author: rana 
Type: Update 
Category: LSC 
Subject: Moved LSC sensing matrix notch frequencies 

We looked at the DRMI noise spectrum and chose new excitation frequencies such that the lines are lower in frequency than before (~300 Hz instead of 800 Hz) and also not in some noisy region.

New filters is saved and loaded for all LSC DOFs.

Attachment 1: NewLSCnotches.png  64 kB  | Hide | Hide all
NewLSCnotches.png
ELOG V3.1.3-